The sum of three consecutive odd integers is 39. Write equations to model the situation, and find the values of the three integers
Answer by TimothyLamb(4379)

z = y + 2
y = x + 2
x + y + z = 39
---
put the system of linear equations into standard form
---
y - z = -2
x - y = -2
x + y + z = 39
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 11
y = 13
z = 15
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---